A balloon is inflated with helium to a volume of 0.75 L at 27°C. The balloon was found several  hours later, snd the temperature had dropped to 22°C. What would be the volume of the balloon when found, if no helium had escaped? 



1
Expert's answer
2021-03-08T06:04:32-0500

V1 = 0.75L , T1 = 27⁰C ,V2 = ? , T2 = 22⁰C

V1/T1 = V2/T2

0.75/27 = V2/22

V2 = (0.75×22)/27

V2 = 0.61 L